OpenStudy (anonymous):

Julie is solving the equation x^2 + 6x + 9 = 0 and notices that the discriminant b^2 - 4ac has a value of 0. This tells her that the equation has A) no real roots. B) one real root. Eliminate C) two real roots. D) three real roots.

OpenStudy (accessdenied):

If the discriminant is 0, look at the quadratic formula where the discriminant is 0 \[ x = \frac{-b \pm \sqrt{\color{goldenrod}{b^2 - 4ac}}}{2a} \]

OpenStudy (accessdenied):

If that part is 0, then the square root just cancels out, and whether you add or subtract 0, you get the same number. So... \[ x = -\! \frac{b}{2a} \]

OpenStudy (anonymous):

so it would be A right

OpenStudy (anonymous):

no real roots

OpenStudy (accessdenied):

Nope, x=-b/2a is a real value. Just in general terms. You have one real root, which has multiplicity 2.

OpenStudy (accessdenied):

You could also factor it to figure it out, its a perfect square trinomial that you can factor into (x+3)^2 x=-3